OpenStudy (anonymous):

Evaluate the difference quotient for the given function. f(x) = (x+5)/(x + 2) (f(x) − f(1))/(x − 1)

OpenStudy (anonymous):

ick!

OpenStudy (anonymous):

lets make life easy and compute \[f(1)=\frac{1+5}{1+2}=\frac{6}{3}=2\] forst

OpenStudy (anonymous):

so you have \[\frac{f(x)-2}{x-1}\] as a start

OpenStudy (anonymous):

i should just quit. i keep putting x + 1 when it is x + 2 this time for real \[f(x)-2=\frac{x+5}{x+2}-2=\frac{x+5-2(x+2)}{x+2}\] \[=\frac{x+5-2x-4}{x+2}=\frac{1-x}{x+2}\] there now it is right!

OpenStudy (anonymous):

now divide by \[x-1\] and get \[\frac{1-x}{(x-1)(x+2)}=\frac{-1}{x+2}\]

OpenStudy (anonymous):

because \[\frac{1-x}{x-1}=-1\]

OpenStudy (anonymous):

sorry i messed up twice with those stupid typos. it is right now

OpenStudy (anonymous):

thank you for the step-by-step analysis!